Janine is 25 and has a good job at a biotechnology company. Janine estimates that she will need $875,000 in her total retirement nest egg by the time she is 65 in order to have retirement income of $20,000 a year. (She expects that Social Security will pay her an additional $15,000 a year.) She currently has $5,000 in an IRA, an important part of her retirement nest egg. She believes her IRA will grow at an annual rate of 8 percent, and she plans to leave it untouched until she retires at age 65. How much money will Janine have to accumulate in her company’s 401(k) plan over the next 40 years in order to reach her retirement income goal?

401(k) accumulation ?
Answer #1

IRA Amount = $ 5000 and Rate of Growth of IRA = 8 % , Tenure of Growth = 40 years

IRA Amount at Retirement = 5000 x (1.08)^(40) = $ 108622.6

Target Retirement Nest Egg = $ 875000

Therefore, Required Accumulation in 401(k) = 875000 - 108622.6 = $ 766377.4